### v1

@@ Abstract
In the advancement of industrial informatization, Unsupervised Industrial Anomaly Detection (UIAD) technology effectively overcomes the scarcity of abnormal samples and significantly enhances the automation and reliability of smart manufacturing. While RGB, 3D, and multimodal anomaly detection have demonstrated comprehensive and robust capabilities within the industrial informatization sector, existing reviews on industrial anomaly detection have not sufficiently classified and discussed methods in 3D and multimodal settings. We focus on 3D UIAD and multimodal UIAD, providing a comprehensive summary of unsupervised industrial anomaly detection in three modal settings. Firstly, we compare our surveys with recent works, introducing commonly used datasets, evaluation metrics, and the definitions of anomaly detection problems. Secondly, we summarize five research paradigms in RGB, 3D and multimodal UIAD and three emerging industrial manufacturing optimization directions in RGB UIAD, and review three multimodal feature fusion strategies in multimodal settings. Finally, we outline the primary challenges currently faced by UIAD in three modal settings, and offer insights into future development directions, aiming to provide researchers with a thorough reference and offer new perspectives for the advancement of industrial informatization. Corresponding resources are available at https://github.com/Sunny5250/Awesome-Multi-Setting-UIAD.
@@ 1 Introduction
Although the field of multimodal anomaly detection has made significant progress in recent years, especially driven by deep learning and sensor technologies, a comprehensive overview summarizing the progress and its practical applications in industrial systems is currently not available. Table 1 shows the advantages of our survey compared with previous surveys of industrial anomaly detection. We provide a more comprehensive summary of methods and datasets in the fields of pure RGB, pure 3D, and multimodal anomaly detection, and discuss relevant research directions based on the needs of the industrial manufacturing. To the best of our knowledge, this survey is the first to systematically classify and deeply review UIAD in different settings, with special attention to the starting point of their design and the challenges they face when applied in practice. The framework roadmap of this survey is shown in Fig. 3. The paper mainly makes the following contributions:
  • To the best of our knowledge, we are the first to present a comprehensive survey for the unsupervised industrial anomaly detection task in different settings (RGB, 3D and multimodal).
@@
2 Datasets, Metrics and Methods
@@ 2.2.3 Instance-Level Metrics
2.3 Problem Definition
Unsupervised anomaly detection refers to using unsupervised methods to detect and localize anomaly patterns or outliers in samples through one or more modals of data. First, we use a training set $\mathcal{X}=\left\{x_{1},x_{2},\ldots,x_{N}\right\}$ containing only normal data samples to train an anomaly detection model $f$ that learns representations of normal patterns. The training goal is to find the optimal model parameters so that the model $f$ can effectively represent the distribution of normal data, which can be expressed by the following optimization problem:
$s_{test}=f\left(x_{test};\theta^{*}\right)$ (2)
@@ 5.3 Multimodal Feature Fusion
5.3.1 Early Fusion
The process of early fusion is shown in Fig. 13. Early fusion refers to the fusion of different modal features before entering the main detection network (for example, after the feature extraction stage), with subsequent networks receiving only a single fused feature as input. Early fusion eliminates the need to have separate branches for different modals in the detection network, reduces the network parameters and training costs. However, the disadvantage is the difficulty in handling the overwhelming influence of a single modal, which can lead to information imbalance. AST [103] directly concatenates RGB features and depth maps along the channel dimension as fused features before feeding them into the teacher-student network, and serves the fused features as the inputs of the subsequent network. ITNM [121] uses pre-trained feature extractors to extract RGB features and point cloud features, and concatenates the weighted point cloud features with RGB features.
@@ 5.3.1 Early Fusion
5.3.2 Middle Fusion
The process of middle fusion is shown in Fig. 14. Middle fusion refers to the fusion of different modal features in the subject detection network and uses the fused features as the inputs of the subsequent network. Compared to early fusion, middle fusion retains separate processing networks for different modal features, and uses additional processing networks to fuse multimodal information and make decisions to obtain anomaly detection results. EasyNet [24] does not use pre-trained feature extractors, but directly inputs RGB images and depth maps into reconstruction networks separately. EasyNet uses the features of the reconstruction process of different modals as the input of the subsequent fusion module, and obtains the anomaly detection result after an attention-based information entropy fusion module and a discriminator network. 3DSR [142] employs channel-separated multimodal reconstruction network without multimodal feature interaction to reconstruct RGB images and depth maps respectively, and concatenates the RGB images and depth maps before and after reconstruction, feeding them into a subsequent anomaly detection module to obtain detection results. DBRN [14] is similar to EasyNet. DBRN uses a dual-branch reconstruction network to reconstruct the input RGB images and depth maps respectively, and then concatenates the RGB images and depth maps before and after reconstruction, feeding them into a subsequent discriminator network.
@@ 5.3.3 Late Fusion
The process of late fusion is shown in Fig. 15. Late fusion refers to the fusion of detection results of different modal networks after the main detection network. Late fusion discards the interaction between multimodal features, retains the independence of information from different modals and avoids information interference between modals. However, this strategy increases the model parameters and complexity, as it requires designing separate processing networks for different modals. 3DRÆM [143] extends DRÆM [139] to multimodal inputs without changing the network architecture. 3DRÆM uses the same reconstruction and discrimination networks for both RGB images and depth maps, and aggregates the final anomaly detection results to obtain the final detection results.
5.3.4 Hybrid Fusion
@@
Hybrid fusion is a combination of the above strategies, which applies different fusion strategy at different processing stages to fully interact features across different modals, exploiting the complementary advantages of multimodal information. Compared to the above three types of fusion strategies, hybrid fusion has greater flexibility and is applicable to different types of models. Hybrid fusion is also the most commonly used multimodal feature fusion strategy. M3DM [123] does not directly concatenate different modal features like previous methods. Instead, M3DM aligns 3D point cloud features to the 2D plane in the feature extraction stage, and uses contrastive learning to fuse the two modal features, encouraging the same position of different modal features to have more corresponding information. M3DM also utilizes a 3D feature memory bank, a RGB feature memory bank, and a fused feature memory bank to make decisions. CPMF [19] presents point clouds as multi-view 2D images during the feature extraction stage and extracts point cloud features and 2D features. Then CPMF concatenated the two modal features as fusion features, which are used to update the memory bank and for inference. Shape-Guided [29] uses pre-trained feature extractors to separately extract RGB features and point cloud features, utilizing point cloud patch features to guide the retrieval of RGB patch features. Shape-Guided constructs a dual-memory bank to reconstruct RGB features and point cloud features. Finally, the anomaly detection results are obtained through the combined reconstruction loss of the two modals. CFM [30] projects point cloud features onto a 2D plane to obtain 3D features. CFM extracts the relationships between features from different modal information through cross-modal mapping, and aggregates the reconstruction errors of the two modal features as anomaly detection results. LSFA [116], based on M3DM, uses contrastive learning to fuse multimodal features from local and global views. LSFA uses a RGB memory bank and a point cloud memory bank to infer the anomaly detection results respectively. The final detection results is obtained by aggregating the multimodal detection results. MMRD [45] uses parameter-free modal modulation in a frozen teacher network encoder to realize multimodal feature interaction and obtain fused features. MMRD then performs intra-modal interaction and inter-modal interaction in a student network decoder to obtain new fused features, and calculates pixel-wise similarity with the fused features obtained from the teacher network encoder to obtain anomaly detection results. CMDIAD [111] first trains a cross-modal distillation network to establish mappings from the feature space of one modal to the feature space of another modal, then uses memory banks to infer anomaly detection results separately for RGB features and point cloud features, and uses OCSVM to implement decision layer fusion.
@@ 7.3 Multimodal UIAD
1) Difficult to solve the problem of missing modals and data noise. Most of the existing multimodal UIAD methods use RGB images and 3D information as input, with both modals jointly providing criteria for anomaly detection. In actual industrial production lines, data from some modals may be missing or contain noise.
@@ 8.3 Multimodal UIAD
Optimize algorithms to compensate for information gaps caused by modal missing or noise, enhancing denoising and anti-interference capabilities. Research algorithms capable of aligning different modal data and effectively achieving multimodal information fusion to address the issue of incomplete alignment of multimodal data. In multimodal detection methods, consider optimization directions such as lightweight models, real-time inference, multi-class detection, few-shot detection, and zero-shot detection to improve deployment effect on actual production lines.
@@ 9 Conclusion
In this paper, we provide a comprehensive review of multimodal UIAD methods and multimodal anomaly detection datasets. Through comprehensive analysis, we reveal the strengths and limitations of existing methods and datasets. Our analysis highlights the potential of multimodal anomaly detection in improving detection accuracy and reducing inference cost while pointing out the challenges and future directions for the field. The contribution of this analysis is to systematically classify and review various methods and datasets for multimodal UIAD for the first time, and analyze their design starting points and challenges in practical applications, providing valuable references and directions for future research. We hope that this paper can provide the latest information reference for the research on multimodal anomaly detection tasks and provide suggestions for future development directions.

### v2

@@ Abstract
In the advancement of industrial informatization, unsupervised anomaly detection technology effectively overcomes the scarcity of abnormal samples and significantly enhances the automation and reliability of smart manufacturing. As an important branch, industrial image anomaly detection focuses on automatically identifying visual anomalies in industrial scenarios (such as product surface defects, assembly errors, and equipment appearance anomalies) through computer vision techniques. With the rapid development of Unsupervised industrial Image Anomaly Detection (UIAD), excellent detection performance has been achieved not only in RGB setting but also in 3D and multimodal (RGB and 3D) settings. However, existing surveys primarily focus on UIAD tasks in RGB setting, with little discussion in 3D and multimodal settings. To address this gap, this artical provides a comprehensive review of UIAD tasks in the three modal settings. Specifically, we first introduce the task concept and process of UIAD. We then overview the research on UIAD in three modal settings (RGB, 3D, and multimodal), including datasets and methods, and review multimodal feature fusion strategies in multimodal setting. Finally, we summarize the main challenges faced by UIAD tasks in the three modal settings, and offer insights into future development directions, aiming to provide researchers with a comprehensive reference and offer new perspectives for the advancement of industrial informatization. Corresponding resources are available at https://github.com/Sunny5250/Awesome-Multi-Setting-UIAD.
@@ 1 Introduction
Industrial anomaly detection involves various modal inputs, such as images, videos, time series, etc., with different modal inputs corresponding to different tasks. We focus on image anomaly detection, which is one of the widely discussed tasks in computer vision. Unlike tasks such as video and time series anomaly detection, image anomaly detection detects anomalies on the surface or structure of objects by extracting visual features from images. With the increasing maturity of RGB and 3D imaging technologies, industrial vision detection systems are gradually transitioning from single-modal to multimodal. RGB images, due to their rich color and texture information, remain the primary data source in most industrial applications. However, 3D data (such as point clouds and depth maps) can provide additional spatial information, enhancing the accuracy and robustness of detection. Moreover, multimodal approaches combine different types of visual modal data (such as combining RGB and 3D point clouds), enabling more comprehensive detection of various types of industrial anomalies and addressing potential information gaps in single-modal approaches.
@@
Although the field of multimodal anomaly detection has made significant progress in recent years, especially driven by deep learning and sensor technologies, a comprehensive overview summarizing the progress and its practical applications in industrial systems is currently not available. Table 1 shows the advantages of our survey compared with previous surveys of industrial anomaly detection. We provide a more comprehensive summary of methods and datasets in the fields of pure RGB, pure 3D, and multimodal anomaly detection, and discuss relevant research directions based on the needs of the industrial manufacturing. From the perspective of modal settings, we introduce detection tasks in different modal settings to meet the varying needs of different production lines for different types of data. For example, production lines for low-cost, simple products may only require RGB images to complete the detection process. In contrast, production lines for high-cost, complex products may need to combine high-resolution RGB images with 3D point cloud data to complete the detection process. To the best of our knowledge, this survey is the first to systematically classify and deeply review UIAD in different settings, with special attention to the starting point of their design and the challenges they face when applied in practice. The framework roadmap of this survey is shown in Fig. 3. The paper mainly makes the following contributions:
  • To the best of our knowledge, we are the first to present a comprehensive survey for the unsupervised industrial image anomaly detection task in different settings (RGB, 3D and multimodal).
@@
2 Preliminaries
2.1 Unsupervised Anomaly Detection
Unsupervised anomaly detection refers to using unsupervised methods to detect and localize anomaly patterns or outliers in samples through one or more modals of data. The unsupervised industrial image anomaly detection discussed in this artical uses normal sample images as training data, allowing the model to learn the features of the normal pattern in the images. First, we use a training set $\mathcal{X}=\left\{x_{1},x_{2},\ldots,x_{N}\right\}$ containing only normal data samples to train an anomaly detection model $f$ that learns representations of normal patterns. The training goal is to find the optimal model parameters so that the model $f$ can effectively represent the distribution of normal data, which can be expressed by the following optimization problem:
$s_{test}=f\left(x_{test};\theta^{*}\right)$ (2)
@@ 5.3 Multimodal Feature Fusion
5.3.1 Fusion Strategies
Early Fusion. The process of early fusion is shown in Fig. 13. Early fusion refers to the fusion of different modal features before entering the main detection network (for example, after the feature extraction stage), with subsequent networks receiving only a single fused feature as input. Early fusion eliminates the need to have separate branches for different modals in the detection network, reduces the network parameters and training costs. However, the disadvantage is the difficulty in handling the overwhelming influence of a single modal, which can lead to information imbalance. AST [105] directly concatenates RGB features and depth maps along the channel dimension as fused features before feeding them into the teacher-student network, and serves the fused features as the inputs of the subsequent network. ITNM [123] uses pre-trained feature extractors to extract RGB features and point cloud features, and concatenates the weighted point cloud features with RGB features.
@@ 5.3.1 Fusion Strategies
Middle Fusion. The process of middle fusion is shown in Fig. 14. Middle fusion refers to the fusion of different modal features in the subject detection network and uses the fused features as the inputs of the subsequent network. Compared to early fusion, middle fusion retains separate processing networks for different modal features, and uses additional processing networks to fuse multimodal information and make decisions to obtain anomaly detection results. EasyNet [24] does not use pre-trained feature extractors, but directly inputs RGB images and depth maps into reconstruction networks separately. EasyNet uses the features of the reconstruction process of different modals as the input of the subsequent fusion module, and obtains the anomaly detection result after an attention-based information entropy fusion module and a discriminator network. 3DSR [144] employs channel-separated multimodal reconstruction network without multimodal feature interaction to reconstruct RGB images and depth maps respectively, and concatenates the RGB images and depth maps before and after reconstruction, feeding them into a subsequent anomaly detection module to obtain detection results. DBRN [14] is similar to EasyNet. DBRN uses a dual-branch reconstruction network to reconstruct the input RGB images and depth maps respectively, and then concatenates the RGB images and depth maps before and after reconstruction, feeding them into a subsequent discriminator network.
@@
Late Fusion. The process of late fusion is shown in Fig. 15. Late fusion refers to the fusion of detection results of different modal networks after the main detection network. Late fusion discards the interaction between multimodal features, retains the independence of information from different modals and avoids information interference between modals. However, this strategy increases the model parameters and complexity, as it requires designing separate processing networks for different modals. 3DRÆM [145] extends DRÆM [141] to multimodal inputs without changing the network architecture. 3DRÆM uses the same reconstruction and discrimination networks for both RGB images and depth maps, and aggregates the final anomaly detection results to obtain the final detection results.
@@
Hybrid Fusion. Hybrid fusion is a combination of the above strategies, which applies different fusion strategy at different processing stages to fully interact features across different modals, exploiting the complementary advantages of multimodal information. Compared to the above three types of fusion strategies, hybrid fusion has greater flexibility and is applicable to different types of models. Hybrid fusion is also the most commonly used multimodal feature fusion strategy. M3DM [125] does not directly concatenate different modal features like previous methods. Instead, M3DM aligns 3D point cloud features to the 2D plane in the feature extraction stage, and uses contrastive learning to fuse the two modal features, encouraging the same position of different modal features to have more corresponding information. M3DM also utilizes a 3D feature memory bank, a RGB feature memory bank, and a fused feature memory bank to make decisions. CPMF [19] presents point clouds as multi-view 2D images during the feature extraction stage and extracts point cloud features and 2D features. Then CPMF concatenated the two modal features as fusion features, which are used to update the memory bank and for inference. Shape-Guided [29] uses pre-trained feature extractors to separately extract RGB features and point cloud features, utilizing point cloud patch features to guide the retrieval of RGB patch features. Shape-Guided constructs a dual-memory bank to reconstruct RGB features and point cloud features. Finally, the anomaly detection results are obtained through the combined reconstruction loss of the two modals. CFM [30] projects point cloud features onto a 2D plane to obtain 3D features. CFM extracts the relationships between features from different modal information through cross-modal mapping, and aggregates the reconstruction errors of the two modal features as anomaly detection results. LSFA [118], based on M3DM, uses contrastive learning to fuse multimodal features from local and global views. LSFA uses a RGB memory bank and a point cloud memory bank to infer the anomaly detection results respectively. The final detection results is obtained by aggregating the multimodal detection results. MMRD [45] uses parameter-free modal modulation in a frozen teacher network encoder to realize multimodal feature interaction and obtain fused features. MMRD then performs intra-modal interaction and inter-modal interaction in a student network decoder to obtain new fused features, and calculates pixel-wise similarity with the fused features obtained from the teacher network encoder to obtain anomaly detection results. CMDIAD [113] first trains a cross-modal distillation network to establish mappings from the feature space of one modal to the feature space of another modal, then uses memory banks to infer anomaly detection results separately for RGB features and point cloud features, and uses OCSVM to implement decision layer fusion.
5.3.2 Fusion Methods
Existing multimodal feature fusion methods can be divided into three categories: concatenation-based methods, projection-based methods, and learning-based methods. (1) Concatenation-based methods are the most straightforward fusion approach, where different modal features are directly concatenated along the channel dimension to obtain fused features, which are then used as input for subsequent models. (2) Projection-based methods map different modal information into a shared feature space, achieving information fusion between modals. These methods can also be combined with other fusion operations, such as concatenation after projection into the shared feature space. (3) Learning-based methods use deep learning or contrastive learning to learn how to effectively fuse features from different modals. These methods can better capture the complex relationships between modals but also face additional challenges, such as increased training costs and the risk of underfitting or overfitting.
Concatenation-based methods are often used to fuse RGB feature maps and depth maps. When using 3D point cloud features instead of depth maps, downsampling is typically required to match the size of the RGB feature map. AST [105] downsamples both RGB image features and depth maps to the same size and concatenates them along the channel dimension to obtain the fused features. DBRN [14], 3DRÆM [145], and 3DSR [144] are similar to AST, directly concatenating RGB image features and depth maps along the channel dimension and feeding them into subsequent anomaly detection (discriminative) modules. ITNM [123] considers that the feature vectors of the two modals differ in both dimension and scale, so it applies weighting to the downsampled point cloud features to balance them before concatenating with RGB features. CPMF [19] normalizes the RGB features and downsampled point cloud features before concatenating them along the channel dimension. Shape-Guided [29] applies an affine transformation to the RGB anomaly scores to align them with the 3D scores, then directly fuses the two modals’ anomaly scores by taking the maximum score from each pixel. EasyNet [24] calculates the information entropy of channels with only RGB features and the information entropy of channels with both RGB and depth features before concatenating the multimodal features. If the fusion of depth features enhances the information entropy, the features of both modals are then concatenated.
Projection-based methods are commonly used to fuse RGB features and 3D point cloud features. M3DM [125] projects the extracted 3D point cloud features onto a 2D plane using point coordinates and camera parameters, and unifies the sizes of 2D patch features and 3D patch features through average pooling. Similar to M3DM, CFM [30] and LSFA [118] interpolate and project 3D point cloud features onto a 2D plane, thus enabling further fusion of features from the two modals.
Learning-based methods, compared to the previous two methods, have a broader application domain. M3DM proposes using contrastive learning to fuse RGB features and projected 3D point cloud features. It encourages multimodal patch features at the same position to have the most mutual information through patch-wise contrastive loss, meaning that the diagonal elements of the contrastive matrix have the highest values. LSFA, inspired by M3DM, uses contrastive learning to fuse multimodal features while adding an interaction of global structural information. It clusters local features to obtain instance-wise features, and then applies contrastive loss to the instance-wise features. CMDIAD [113] does not focus on information fusion between the two modals but instead establishes a mapping from the feature space of one modal to the feature space of the other modal. This results in a cross-modal distillation network that enables interaction between different modals.
5.3.3 Fusion Strategy Optimization
The four multimodal feature fusion strategies introduced in Section 5.3.1 are the mainstream approaches currently used in multimodal industrial anomaly detection tasks. However, considering the high complexity of industrial scenarios and the high requirements for detection accuracy, there is still room for optimization in the existing strategies.
In terms of fusion efficiency, early fusion can reduce training costs by fusing features before detection, but it struggles to handle the overwhelming influence of a single modal, which may lead to information imbalance. Middle fusion fuses features at an intermediate stage, allowing better retention of the characteristics of each modal, but determining the optimal fusion timing is challenging and may affect feature interaction, then impacting the detection results. Late fusion combines results after detection, avoiding information interference, but it increases model parameters and processing complexity. Hybrid fusion is highly flexible, but in practical applications, the challenge remains to precisely select the fusion stage and method for different industrial scenarios, which is still an unresolved issue.
Moreover, the current strategies are difficult to adapt to certain anomaly situations in complex and dynamic industrial environments, such as missing modals. In multimodal industrial anomaly detection tasks, missing modal data may occur due to factors such as equipment failures and environmental interference. Existing multimodal detection methods require joint input from all modals, and the absence of one modal can prevent the detection process from proceeding.
3D-ADNAS [86] optimizes multimodal anomaly detection methods from the perspective of mult-modal fusion architecture design. It introduces Neural Architecture Search (NAS), designing a two-level search space that simultaneously optimizes both intra-module (feature selection and fusion within modules) and inter-module (searching for the optimal module combination strategy) aspects. By combining gradient optimization strategies, it automatically searches for an efficient multimodal fusion architecture. 3D-ADNAS expands the perspective of multimodal fusion beyond the four previously mentioned strategies, combining existing strategies in a more flexible way to obtain more effective solutions.
RADAR [92] focuses on the problem of modal-missing in multimodal industrial anomaly detection. It introduces modality-incomplete instruction, using a pre-trained multimodal Transformer to dynamically adapt to missing modals. RADAR designs a real-pseudo label hybrid module to highlight the uniqueness of different modal combinations, enhancing anomaly localization ability, and combines a multi-feature repository (RGB, point clouds and fused features) to compute anomaly scores. RADAR effectively mitigates the impact of modal-missing cases on multimodal industrial anomaly detection tasks and presents this issue as a new task.
@@ 7.3 Multimodal UIAD
1) Difficult to solve the problem of missing modals and data noise. Existing multimodal industrial anomaly detection methods require all modals to provide criteria jointly. In actual industrial production lines, due to sensor failures or environmental factors, data from some modals may be missing or contain noise. Additionally, considering the complexity of real-world production lines, other modals beyond RGB images and 3D point clouds may also be required, such as infrared images, grayscale images, and other monochrome images.
@@ 8.3 Multimodal UIAD
Optimize algorithms to compensate for information gaps caused by modal missing or noise, enhancing denoising and anti-interference capabilities. Consider incorporating more modals into the dataset and methods, such as infrared images, grayscale images, and others, to create a unified dataset that includes multiple modals such as RGB, 3D, and infrared. Methods based on this dataset can select specific modals for anomaly detection. Research algorithms capable of aligning different modal data and effectively achieving multimodal information fusion to address the issue of incomplete alignment of multimodal data. In multimodal detection methods, consider optimization directions such as lightweight models, real-time inference, multi-class detection, few-shot detection, and zero-shot detection to improve deployment effect on actual production lines.
@@ 9 Conclusion
In this paper, we provide a comprehensive review of the latest progress in unsupervised industrial image anomaly detection in RGB, 3D, and multimodal settings. Based on the concept, datasets, and evaluation metrics of unsupervised industrial image anomaly detection, we review and summarize the research from the perspective of different modal settings, and classify the multimodal feature fusion strategies in multimodal setting. Additionally, considering the algorithm deployment in practical production lines, we summarize methods such as few-shot learning, zero-shot learning, multi-class learning, and anomaly simulation learning. Through a comprehensive analysis, we highlight the challenges faced by anomaly detection research in different modal settings in real-world applications, providing valuable references and directions for future research. We hope that this paper can provide the latest information reference for the research on multimodal anomaly detection tasks and provide suggestions for future development directions.